# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### v2

@@ D.1 Domain adaptation methods
| Model | mAP {10,30,50}% | Avg | ||
| ADDA [40] | 30.16 | 27.58 | 19.72 | 25.82 |
| WDGRL [42] | 24.77 | 21.96 | 14.81 | 20.51 |
| DANN [1] | 31.32 | 28.55 | 20.10 | 26.65 |
| MSTN [2] | 30.17 | 27.32 | 18.55 | 25.34 |
| SSTDA [7] | 29.58 | 26.81 | 18.36 | 24.92 |
| TranSVAE [65] | 28.31 | 25.67 | 16.72 | 23.57 |
| Ours (SADA) | 31.54 | 28.77 | 20.22 | 26.84 |
D.2 Ablation over variants of the SADA loss
@@
| Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
| 34.41 | 30.58 | 21.00 | 28.66 | |||
| 34.11 | 30.48 | 20.75 | 28.44 | |||
| 34.46 | 30.85 | 20.90 | 28.73 | |||
| 34.63 | 31.08 | 21.26 | 28.99 | |||
| 33.47 | 29.91 | 20.17 | 27.85 | |||
| 32.84 | 29.06 | 19.41 | 27.10 | |||
| 34.86 | 31.16 | 21.46 | 29.16 | |||
[…]
| Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
| 29.65 | 26.86 | 19.14 | 25.21 | |||
| 31.32 | 28.55 | 20.10 | 26.65 | |||
| 29.94 | 27.39 | 19.40 | 25.57 | |||
| 30.86 | 28.02 | 19.84 | 26.57 | |||
| 30.84 | 27.98 | 19.37 | 26.06 | |||
| 29.77 | 27.10 | 18.70 | 25.19 | |||
| 31.54 | 28.77 | 20.22 | 26.84 | |||
Closely related to Sec. IV.C of the main text, in Table S6 we ablate over different variants of our loss on scenario S2 and S4. These scenarios define the source domain as white old and new counter kitchens and the rest as the target, respectively. Observe in this table that our main insights from Sec. IV.C of the main text are mostly confirmed. Concretely, one of our main observations was that aligning the complete global distribution was generally preferable over aligning the partial background distribution only. This observation is satisfied in most of the scenarios that we study –with the exception of scenario S3. Similarly, in Sec. IV.C of the main text we argued that combining alignment loss that presented overlaps yielded also a decrease in performance. More in detail, this happens when we couple either the local or the background loss with a global alignment approach. In these cases, the embeddings have a double (simultaneous) alignment objective, which we argue is not desirable. As observed in Table S6, this observation is also mostly satisfied in these scenarios as the combination of these losses yields a performance degradation in 7 of the 8 cases studied in this paper.

### v3

@@ H.1 Masking out the anchors during inference
Ours(SADA) | 31.60 | 28.22 | 18.98 | 26.26 | - | |
@@
S1 | ||||||
Filtered % | mAP {10,20,30,40,50}% | Avg % | ||||
0% | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.71 |
25% | 32.76 | 31.60 | 29.81 | 26.14 | 20.24 | 28.11 |
50% | 35.25 | 34.25 | 32.68 | 29.25 | 23.01 | 30.89 |
75% | 38.39 | 37.77 | 36.40 | 33.19 | 26.33 | 34.42 |
100% | 41.68 | 40.59 | 40.23 | 34.26 | 30.03 | 37.35 |
S3 | ||||||
Masked % | mAP {10,20,30,40,50}% | Avg % | ||||
0% | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
25% | 32.54 | 31.87 | 30.06 | 26.99 | 21.91 | 28.67 |
50% | 34.46 | 33.55 | 31.29 | 27.59 | 21.18 | 29.61 |
75% | 35.09 | 34.53 | 32.97 | 30.05 | 23.97 | 31.32 |
100% | 35.70 | 35.24 | 34.09 | 30.72 | 24.15 | 31.98 |
Nevertheless, we highlight that in this scenario in particular, unlike in S3, there is still a big gap which indicates that there is considerable room for improvement. To gain further insights on this issue, we additionally propose an experiment where we progressively mask out the set of background anchors. For this, in Tab. L we randomly mask out these ensuring that only a percentage of the background anchors are not considered during inference. As observed in Tab. L, in both considered scenarios we observe a linear performance increase, reaching an absolute improvement of mAP and mAP of the version with filtered background anchors over the version. This is, the variant where only the action anchors are considered over that containing both action and all the background ones.
H.2 Masking out the anchors during training and inference
In Tab. M we propose an alternative analysis. In this case, we do not only mask the background anchors during inference but also during training. In this case, for instance, masking out of the background anchors additionally implies not training any of the task losses on them, nor using them on our alignment loss. As observed in the results, masking them out entirely – i.e., – has an even bigger impact, reaching an absolute improvement of mAP and mAP, respectively, when comparing it to the filtering version. Nevertheless, in this case, we observe that the improvement is no longer linear, and even worsens the performance for small percentages of masking. We hypothesize this is induced by a reduction of the training data of background class. Thus, while masking some background embeddings has a positive effect, we must also consider the negative effect that having less training data has on the ones that we do consider. Hence, the improvement in performance remains little to non-existent until this trade-off is dominated by the positive effect of masking out the majority of background anchors. In short, these additional experiments emphasize the need for devising carefully designed models that can properly cope with these intrinsic limitations of anchor-based methods.
S1 | ||||||
Filtered % | mAP {10,20,30,40,50}% | Avg | ||||
0% | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.71 |
25% | 32.25 | 30.46 | 28.04 | 24.54 | 19.03 | 26.86 |
50% | 33.43 | 32.06 | 29.98 | 26.44 | 20.55 | 28.49 |
75% | 38.26 | 37.33 | 35.72 | 32.53 | 25.61 | 33.89 |
100% | 47.08 | 46.69 | 45.65 | 42.31 | 34.42 | 43.23 |
S3 | ||||||
Masked % | mAP {10,20,30,40,50}% | Avg | ||||
0% | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
25% | 31.48 | 30.34 | 27.97 | 24.07 | 18.93 | 26.56 |
50% | 32.11 | 31.11 | 29.25 | 25.85 | 20.72 | 27.81 |
75% | 35.99 | 35.28 | 33.48 | 29.90 | 23.64 | 31.66 |
100% | 37.06 | 35.73 | 32.34 | 26.45 | 33.82 | 47.08 |
I Additional qualitative results
In this section we extend our qualitative analysis from Sec. 4.3. To this end, we first provide in Sec. I.1 additional segment visualizations. Then, in Sec. I.2 we extend the analysis of the TSNE plots to study the effect of multi-resolution in the overall alignment of embeddings.
I.1 Segment visualization
In Fig. E we present a segment visualization of S3 – i.e., using old dark-counter kitchens as the source domain, and the rest as the target. Observe that in this case, both ActionFormer [63] and Tridet [48] perform similarly. Concretely, they mainly miss one close action at the beginning, a put in the middle section of the video, and finally the last two actions open and take. In contrast, our model is able to correctly detect the previously missing close, and the final take. In short, reducing by half the number of undetected actions in the video.
Similarly Fig. F shows a segment visualization that compares the chosen baselines with SADA when applied to the CharadesEgo dataset [50]. Observe that this visualization presents a much more complex setup where there are numerous overlapping actions. In this regard, we highlight that the worst-performing model is the ActionFormer [63]. This model misses the action walk and one of the actions put. It also presents important limitations in terms of localization, especially for the action labels drink and hold. The Tridet [48] considerably improves this method. Nevertheless, it still misses the action walk and is unable to locate the longest put action. Additionally, Tridet predicts an action take that does not correspond to any ground truth. Our method, finally, presents much more accurate scores. Importantly, our method does not miss any of the ground-truth actions and in most cases reports considerably accurate segments. We only highlight the persisting confusion in the two actions hold which is consistent across all the tested baselines.
I.2 Extended TSNE analysis: Studying the effect of multi resolution
In the qualitative study presented in the main paper (see Fig. 5) we only include visualizations of the embeddings that correspond to the level 0 of our multi-resolution architecture. It remains unclear, however, whether this alignment behavior remains consistent across the different resolution levels. To showcase this behavior, in this section we first include in Fig. G - Fig. J the TSNE plots of the 3 majority classes and of background embeddings, respectively. Each of these figures includes the different plots at each of the first 3 resolution levels, where level 0 is the shallowest, and level 3 is the deepest studied level. We highlight that deeper levels are not meaningful to plot given the scarce number of resulting domain-invariant embeddings, caused by the downsampling that each of the levels of the architecture performs.
Observe that the influence of the alignment loss follows a similar pattern in all the resolutions. Concretely, in the three presented, we observe a considerable mixing between the source and target distributions (red and blue, respectively). This contrasts with the very clear disentanglement of the representations of both domains in the source-only version, where these present little to no overlap. We highlight what seems to be the only exception which is resolution level 3 of class 1. This one shows little improvement in the mixing of the distributions compared to the source-only variant. We also find that our overall improvement of the mixing of the distributions is consistent when analyzing the background class embeddings. Observe in Fig. J that in all 3 studied resolution levels, SADA improves very considerably the alignment, pushing the feature space of both domains to be domain invariant. This emphasizes the positive influence of the background alignment term of our loss (see Eq. 8).